OpenStudy (anonymous):

Write the equation of a line in point-slope form that has a slope of -1 and passes through the point (-2, 5) A. y + 5 = -1(x – 2) B. y + 2 = 1(x – 5) C. y + 1 = 5 (x + 2) D. y – 5 = -1(x + 2)

OpenStudy (aripotta):

y - y1 = m(x - x1) our point is (-2,5) so -2 would be x1 and 5 would be y1

OpenStudy (aripotta):

and -1 is m. so now just plug those in

OpenStudy (anonymous):

i dont follow

OpenStudy (aripotta):

y - y1 = m(x - x1) is point-slope form. they gave us a point (-2,5) and the slope, -1. all coordinates are setup as (x,y) so x1 would be -2, y1 would be 5, and m would be -1. just substitute those numbers for the corresponding variable, and there you go

OpenStudy (anonymous):

i understand now, i was over thinking it

OpenStudy (aripotta):

oh ok. that's good :)

OpenStudy (anonymous):

thank you so much

OpenStudy (aripotta):

no problemo